OpenStudy (baldymcgee6):

Rotating the function y= e^(-0.5x) about the x-axis on the interval [-1, 0] yields a solid of revolution with a volume of?

OpenStudy (amistre64):

what have you drawn out so far?

OpenStudy (baldymcgee6):

well i dont know how to find the integral of that function

OpenStudy (amistre64):

of what function? what have you worked out so far?

OpenStudy (baldymcgee6):

the function y= e^(-0.5x)... i need\[:\pi \int\limits\limits_{-1}^{0} y^2\]

OpenStudy (amistre64):

looks good so far (e^-.5x)^2 is equal to what?

OpenStudy (amistre64):

this turns out to be a pretty elementary function after all

OpenStudy (baldymcgee6):

e^-1x

OpenStudy (amistre64):

correct

OpenStudy (amistre64):

and if you want to, we can turn this about and get the same results from:\[pi\int_{0}^{1}e^x dx\]

OpenStudy (baldymcgee6):

lets just try and stick with the original equation, dont want to confuse me ;)

OpenStudy (amistre64):

:) k

OpenStudy (baldymcgee6):

so is it like pi(e^0-e^-1)

OpenStudy (amistre64):

almost lets work the integration and see if we get a good setup to play with

OpenStudy (amistre64):

e^-x comes from something akin to e^-x to begin with; but its missing something

OpenStudy (amistre64):

can you do a usubstitution on it? if you really really need to?

OpenStudy (baldymcgee6):

i'm not sure what you mean.. we already have the integral dont we?

OpenStudy (amistre64):

i dont think your setup as correct. id feel better if we stepped thru it to be sure tho

OpenStudy (baldymcgee6):

okay, lets go step by step

OpenStudy (amistre64):

give me your integration of e^-x

OpenStudy (baldymcgee6):

-e^-x

OpenStudy (amistre64):

that looks better; now piece it all in: pi[-e^0-(-e^-1)] pi(-1-(-1/e))

OpenStudy (baldymcgee6):

thats alot of negatives. lol

OpenStudy (amistre64):

:) we can clean it up if need be

OpenStudy (baldymcgee6):

lets do that

OpenStudy (amistre64):

-pi(1-(1/e)) -pi(e-1)/e

OpenStudy (amistre64):

ugh, i think i misplaced a negative along the way

OpenStudy (baldymcgee6):

answers are: A. pi(e+1) B. pi(e^2-1) C. pi(e-1) D. pi(e^2+1)

OpenStudy (amistre64):

-e^-0 - -e^--1 lol

OpenStudy (amistre64):

pi(-1 + e)

OpenStudy (amistre64):

looks like c is our best match

OpenStudy (baldymcgee6):

okay, thats what i think too. thanks for your help. can i ask you another one?

OpenStudy (amistre64):

dunno, i think i used up all my mojo sorting that one out :/

OpenStudy (baldymcgee6):

haha well thanks anyways!

OpenStudy (amistre64):

good luck ;) just post a new question to the left after you close this one out and someones bound to bite